OpenStudy (anonymous):

Find the product of z1 and z2 where z1 = 8(cos 40° + i sin 40°), and z2 = 4(cos 135° + i sin 135°). 32(cos 40° + i sin 40°) 12(cos 175° + i sin 175°) 32(cos 5400° + i sin 5400°) 32(cos 175° + i sin 175°)

OpenStudy (anonymous):

@amistre64 @Nnesha

OpenStudy (amistre64):

theres a general rule, but it seems that you dont know the rule

OpenStudy (amistre64):

multiply it out ....

OpenStudy (anonymous):

like distribute?

OpenStudy (amistre64):

yeah, its basic algebra, with a trig twist :)

OpenStudy (amistre64):

n (cos(a) + i sin(a)) * m (cos(b) + i sin(b)) nm * (cos(a) + i sin(a)) (cos(b) + i sin(b)) the first part of the rule is that "radiuses" multiply so with that we can focus on the trigs

OpenStudy (amistre64):

simplify for the sake of typing (cos(a) + i sin(a)) (cos(b) + i sin(b)) r s p q (r + is)(p + iq)

OpenStudy (anonymous):

32 X (r+ is)(p+iq)

OpenStudy (amistre64):

correct, now expand the right side

OpenStudy (anonymous):

(cos 40° + i sin 40°)(cos 135° + i sin 135°)

OpenStudy (amistre64):

(r+ is)(p+iq) r(p+iq)+ is(p+iq) keep it going

OpenStudy (anonymous):

cos40 (cos 135 + i sin 135)

OpenStudy (amistre64):

we need to keep this as general as possible so that you can see how the stuff works. and not be blind sided by your specific problem values. r(p+iq) + is(p+iq) rp + i rq + i sp + ii sq ^^ ii = -1 rp + i rq + i sp - sq collect like terms, real to real, and i to i (rp - sq) + i (rq + sp) now we can replace them with the general trigs we subbed for, recall: r s p q (cos(a) + i sin(a)) (cos(b) + i sin(b)) real part is: rp - sq imaginary part is: rq + sp real part is: cos(a)cos(b) - sin(a)sin(b) imaginary part is: cos(a)sin(b) + sin(a)cos(b) these trigs are identities for something, do you recall them?

OpenStudy (amistre64):

we are constructing the general rule for these types of problems, so that we dont have to work it out the long way each and every time

OpenStudy (anonymous):

the identities are the sum and difference formulas

OpenStudy (amistre64):

well, the sum formulas yes real part is: cos(a+b) = cos(a)cos(b) - sin(a)sin(b) imaginary part is: sin(a+b) = cos(a)sin(b) + sin(a)cos(b) therefore the rule is the radiuses multiply, the arguments add

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

\[n(\cos(a)+i\sin(a))~*~m(\cos(b)+i\sin(b))\implies nm(\cos(a+b)+i\sin(a+b))\]

OpenStudy (anonymous):

so the 4th option? 32(cos 175° + i sin 175°)

OpenStudy (amistre64):

in a more succinct notation: \[(n,a)(m,b)\implies (nm,a+b)\] \[(8,40)(5,135)\implies (32,175)\]yes

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

also do you know how to do the cube roots? I forgot how to do the degrees

OpenStudy (amistre64):

yeah, i can muddle my way thru it

OpenStudy (anonymous):

do you want a specific example?

OpenStudy (amistre64):

\[(n,a)(n,a)\implies (n^2,2a)\] \[(n,a)(n,a)(n,a)\implies (n^2,2a)(n,a)\implies(n^3,3a)\] as a rule \[(n,a)^k\implies (n^k,ka)\]

OpenStudy (anonymous):

Find the cube roots of 27(cos 279° + i sin 279°).

OpenStudy (amistre64):

cube root is equivalent to ^(1/3)

OpenStudy (anonymous):

yes, but I need to know how to do the degree portion. I know the 3rd root of 27 is 3

OpenStudy (amistre64):

ive already posted the rule ....

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so 3rdrt of 279 and 279/3?

OpenStudy (amistre64):

that is the "primitive" degree for lack of a better term. we add (360/3) to it to find all 3 roots and adjust as needed

OpenStudy (anonymous):

so 279+120 and -120?

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

actually it would be -120 and -240

OpenStudy (amistre64):

279/3 + 120(0) 279/3 + 120(1) 279/3 + 120(2) and adjust as needed to keep them between 0 and 360

OpenStudy (anonymous):

actually it would be -120 and -240

OpenStudy (amistre64):

7th roots would be? 279/7 + 360(0)/7 279/7 + 360(1)/7 279/7 + 360(2)/7 279/7 + 360(3)/7 279/7 + 360(4)/7 279/7 + 360(5)/7 279/7 + 360(6)/7 3^(1/7) of course

OpenStudy (amistre64):

a more convenient way to view this is to divide the circle into its parts, and then add the argument to each part to rotate it appropriately .. does that make sense?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):